1. A piston – cylinder containing steam at 700 kPa and 250 deg C undergoes a constant pressure process until the quality is 70%. Determine per kilogram (a) change of internal energy and (b) change of enthalpy.


1
Expert's answer
2022-01-24T17:09:02-0500

From steam tables at;

"P_1=700kPa" and "T_1=250\u00b0C"

"v_1=0.336m^3\/kg, u_1=2718.66kJ\/kg, h_1=2954.123kJ\/kg"

At 700kPa saturated condition;

"u_f=696.36kJ\/kg,u_g=2971.81kJ\/kg"

"h_f=697.14kJ\/kg,h_g=2762.74kJ\/kg"

"h_2=h_f+x(h_g-h_f)"

x=0.7

"h_2=697.14+0.7(2762.74-697.14)=2143.66kJ\/kg"

"u_2=u_f+x(u_g-u_f)"

"u_2=696.36+0.7(2971.81-696.36)=2009.175kJ\/kg"

a)

"\u2206u=u_2-u_1"

"=2009.175-2718.66"

"-709.485kJ\/kg"

b)

"\u2206h=h_2-h_1"

=2143.66-2954.123

=-810.463kJ/kg